A ball is thrown with an initial velocity of 36 m/s at 33°.



What is the initial velocity of the ball at x ang y-dimension?


What is the final velocity of the ball at x and y- dimension?


What is the time of the ball as it reaches its peak?


What is the maximum height reached of the ball?


What is the total time travel of the ball?

1
Expert's answer
2022-02-23T08:31:01-0500

Given:

"v_0=36\;\rm m/s"

"\theta=33^\circ"

What is the initial velocity of the ball at x ang y-dimension?

"v_{0x}=v_0\cos\theta=30.2\:{\rm m/s}\\
v_{0y}=v_0\sin\theta=19.6\:\rm m/s"

What is the final velocity of the ball at x and y- dimension?

"v_{fx}=v_{0x}=30.2\:{\rm m/s}\\
v_{fy}=-v_{0y}=-19.6\:\rm m/s"

What is the time of the ball as it reaches its peak?

"t=\frac{v_{0y}}{g}=\frac{19.6}{9.8}=2.0\: \rm s"

What is the maximum height reached of the ball?

"h_{\max}=\frac{v_{0y}^2}{2g}=19.6\:\rm m"

What is the total time travel of the ball?

"t_{tot}=2*2.0\: \rm s=4.0\: s"